The tread life of a particular brand of tire is a random variable best described by a normal distribution with a mean of 60,000 miles and a standard deviation of 1600 miles. What is the probability a randomly selected tire of this brand will last between 56,640 miles and 57,120 miles?

i cant seem to figure this out or even know where to begin!... i would be so,so,so greatful if someone would point me in the right direction and at least give me some tools!please!

Answer by stanbon(75887)

The tread life of a particular brand of tire is a random variable best described by a normal distribution with a mean of 60,000 miles and a standard deviation of 1600 miles. What is the probability a randomly selected tire of this brand will last between 56,640 miles and 57,120 miles?
--------------------
Same procedure as on your other problem.
---
Find the z-scores for 56640 and 57120
z(56640)= (56640-60000)/1600 = -2.1
---
z(57120) = (57120-60000)/1600 = -1.8
------------------
P(56640 < x < 57120)= P(-2.1 < z < -1.8) = normalcdf(-2.1,-1.8) = 0.018066..
==================================================================
Cheers,
Stan H.